how do i find both the area and height of an equalateral triangle with all sides 5.4 yeards
Answer by stanbon(75887)

how do i find both the area and height of an equilateral triangle with all sides 5.4 yeards
------
Sketch the triangle.
Draw an altitude from a vertex perpendicular to the opposite side.
---
The altitutde bisects the base.
---
You have 2 right triangles with base = 5.4/2 = 2.7,
height = h, and hypotenuse = 5.4
---
Use Pythagoras to get:
h^2 + 2.7^2 = 5.4^2
h^2 = 5.4^2-2.7^2
height = 4.68
---
Area = (1/2)base*height
= (1/2)5.4*4.68 = 12.63 sq. units
====================================
Cheers,
Stan H.